OpenStudy (anonymous):

a dog breeder wishes to create 5 adjacent pens for his dogs. The breeder has 120ft of fencing to use. What is the largest area he could enclose?

OpenStudy (anonymous):

I have a feeling your question is somewhat incomplete.

OpenStudy (anonymous):

Well the only thing missing is the diagram. I copied the question word for word

OpenStudy (anonymous):

otherwise the area for each pen can be taken to be equal and 120 sqft can be divided by 5. But that is too simple.

OpenStudy (anonymous):

This is calculus, so that seems like it would be way too simple. Also if we did it your way, I think we would be counting each line twice.

OpenStudy (anonymous):

A=6w*l

OpenStudy (anonymous):

Right. w is the width and l is the length. The largest area to be enclosed wrt height is a constant A given by A=6w*l Differenciating with l, A is zero (because it's constant) this 6w+ 6*l*(dw/dl)=0 That gives me dw/dl=-w/l That is the first step.

OpenStudy (anonymous):

Fencing is done around the perimeter, and the perimeter for 5 pens (now, I trust the owner wants equal areas) is 2[(l/5)+w}=P (I trust the owner wants length equally distributed).

OpenStudy (anonymous):

The total perimeter in totality is: the breadth is taken 6 times and length is same, thus P=120ft(given)=2[6w+l] M'kay? Now, differenciate both sides wrt l. That gives you, 6 (dw/dl)+1=0 That is, dw/dl=-1/6 From previous equation, we have dw/dl=-w/l That is, w/l=1/6 That is, l=6w Now 2[6w+l]=120 That is, w=5 and l=30 putting the values you get the area.

OpenStudy (anonymous):

Wow, Thanks for your help

OpenStudy (anonymous):

Just to verify 6(5)*30=900sq ft

OpenStudy (anonymous):

Yes.